# Every SAM Drop Counts: Embracing Semantic Priors for Multi-Modality Image Fusion and Beyond

### v1

@@ Abstract
Multi-modality image fusion, particularly infrared and visible image fusion, plays a crucial role in integrating diverse modalities to enhance scene understanding. Early research primarily focused on visual quality, yet challenges remain in preserving fine details, making it difficult to adapt to subsequent tasks. Recent approaches have shifted towards task-specific design, but struggle to achieve the “The Best of Both Worlds” due to inconsistent optimization goals. To address these issues, we propose a novel method that leverages the semantic knowledge from the Segment Anything Model (SAM) to Grow the quality of fusion results and Establish downstream task adaptability, namely SAGE. Specifically, we design a Semantic Persistent Attention (SPA) Module that efficiently maintains source information via the persistent repository while extracting high-level semantic priors from SAM. More importantly, to eliminate the impractical dependence on SAM during inference, we introduce a bi-level optimization-driven distillation mechanism with triplet losses, which allow the student network to effectively extract knowledge at the feature, pixel, and contrastive semantic levels, thereby removing reliance on the cumbersome SAM model. Extensive experiments show that our method achieves a balance between high-quality visual results and downstream task adaptability while maintaining practical deployment efficiency.
@@ 1 Introduction
Figure 1: Differences between the proposed method and existing mainstream comparative approaches: (a) Traditional and early DL-based methods (e.g., U2Fusion) focus on the fusion visual effect. (b) Task-specific methods (e.g., TarDAL & SegMiF) introduce task loss and features that lead to inconsistent optimization goals, causing a conflict between visual and task accuracy. (c) Our pipeline distills the semantic priors of SAM into a student network, achieving practical inference feasibility while ensuring “the best of both worlds” by SAM’s inherent adaptability to these tasks.
@@
In light of recent advancements in sensing technologies, multi-modality imaging has gained significant attention, with applications in robotics [2], remote sensing [26, 25, 14], and autonomous driving [16, 3]. Integrating infrared and visible sensors provides substantial advantages for intelligent processing [18], although both exhibit inherent limitations [24]. Infrared images are robust against smoke, obstruction, and low light, while visible images excel in resolution, contrast, and texture detail. These characteristics are highly complementary, motivating their fusion to generate a comprehensive fused image that balances visual quality and downstream task requirements. Consequently, achieving this fusion efficiently remains a pressing challenge.
@@
Traditional IVIF methods, based on information theory [17], aim to retain as much source image information as possible but often struggle with optimizing fused image quality, especially in managing redundancy and specific scenarios. Early deep learning-based methods [22, 42, 29, 11, 45, 20] focused on visualization but faced issues like edge blurring and artifacts, failing to meet downstream perception needs. Current methods [10, 12, 33, 21] focus on specific tasks, coupling fusion with downstream processing, resulting in conflicting optimization objectives and challenges in balancing visual quality with task adaptability.
@@
Recent advancements in large-scale visual models have significantly improved various visual analysis tasks [38, 5, 40]. Among these, the Segment Anything Model [8] stands out due to its exceptional and robust capability to provide rich semantic information, making it well-suited for IVIF, as illustrated in Figure 2. Also, its segmentation capabilities align naturally with the requirements of downstream tasks in IVIF field. However, current methods that integrate SAM for low-level tasks typically require full SAM during inference, resulting in excessive practical infeasibility.
@@
Figure 2: Comparative robustness of SAM in MFNet normal scenes (top) and challenging FMB conditions (bottom).
@@
To address all the aforementioned issues, we propose a fusion method named SAGE that fully integrates and distills semantic priors derived from SAM. As illustrated in Figure. 1 III, one of our aim is to harness the advantages of SAM’s semantic priors while reducing computational burden. Specifically, the proposed method consists of two key components: Semantic Persistent Attention (SPA) module and a bi-level distillation scheme. The SPA module focuses on using a persistent repository throughout the process to retain source image information, while heuristically enabling the network to integrate semantic patches predicted by SAM. Furthermore, we introduce a distillation scheme designed to efficiently exclude SAM from the inference process, thereby reducing computational complexity. By enhancing distilled information across semantic, feature, and pixel levels, the triplet losses in the scheme significantly boost the performance of the fusion model. Consequently, efficient fusion is achieved using only the distilled model, delivering high-quality visual results and precise task performance without the direct involvement of SAM. In summary, our contributions can be summarized as follows:
  • We utilize semantic priors from SAM to enhance both visual quality and task accuracy for IVIF.
  • A novel Semantic Persistent Attention (SPA) module is proposed, which efficiently integrates semantic priors while preserving source image information by establishing a persistent repository.
@@
  • We develop a multifaceted optimization distillation scheme that transfers the information processed by SPA into the sub-network, effectively decoupling the fusion process from SAM during inference.
@@ 2 Related Works
Learning based IVIF methods. In recent years, deep learning based multi-modality image fusion approaches [22, 42, 29, 11, 45, 20, 50, 49] achieved significant progress, primarily focusing on improving fusion quality, such as enhancing image details, maintaining structural consistency, and reducing noise. These approaches, including innovations in adversarial learning, Transformer models and diffusion network, have provided new insights. However, most of these methods have not addressed how the fused images can directly contribute to downstream tasks such as object detection, semantic segmentation, and other perception-related applications. With the growing demand for more integrated solutions, this gap has sparked increasing research interest in combining fusion with various downstream tasks.
Early works bridged this gap, with TarDAL [10] introducing a dual-discriminator network for detection optimization, jointly addressing modality differences and detection tasks. Recent methods [36, 37] further focused on detection-optimized fusion. For semantic segmentation [44], SeAFusion [33] and SegMiF [12] improved fusion through task loss and task-specific features, respectively. Recent approaches [13, 21, 32, 19] integrated multi-task learning to jointly optimize detection and segmentation, enhancing performance for both tasks.
@@
SAM & its application.  The Segment Anything Model (SAM) [8] is a large-scale pre-trained foundation model [38] that excels in zero-shot generalization, enabling it to handle a wide range of segmentation and detection tasks without the need for task-specific training. SAM demonstrates impressive performance in low-level vision applications, such as image deblurring [9], dehazing [15], super-resolution [23, 35], and low-light image enhancement [6]. Moreover, SAM’s capabilities extend to high-level tasks like object detection [30], which further strengthens its versatility. This ability to handle both low-level and high-level tasks makes SAM inherently well-suited for the needs of the IVIF field, where fusion based segmentation and detection tasks are critical. However, despite its impressive performance, the high computational complexity presents significant challenges for its practical deployment. Addressing these computational burdens, while fully leveraging semantic priors of SAM, is one of the main primary motivations for our work.
@@ 3.1 Motivation and Problem Formulation
Figure 4: Visual comparison of SOTA approaches across four commonly used datasets, including TNO, RoadScene, M3FD and FMB.
The overall workflow of the proposed method is shown in Figure 3. Our objective is to develop a framework that effectively learns and leverages the semantic priors provided by SAM while minimizing the computational burden during inference. Integrating SAM’s semantic information can significantly enhance the quality of the fused images; however, directly involving SAM make the inference process computationally impractical due to its complexity. To solve this, we utilize a distillation scheme to transfer knowledge from a SAM-integrated main network to a lightweight sub-network. Traditional distillation approaches often face difficulties in capturing the intricate semantic relationships and maintaining the structural consistency necessary for tasks involving cross-modal information fusion. Considering these challenges and the need to jointly optimize both networks for effective semantic transfer, we propose a bi-level optimization framework that models the main and sub-networks as a unified system, aiming to enable a more cohesive and structured transfer of semantic insights.
@@
Denote that $\textbf{I}_{\text{vis}}$ and $\textbf{I}_{\text{ir}}$ represent the visible and infrared input images, respectively. Their corresponding semantic prior patches generated from SAM can be defined as $\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$. The main network, represented by $\mathcal{M}$ with parameters $\boldsymbol{\omega_{\mathrm{m}}}$. while the sub-network, represented by $\mathcal{S}$ with parameters $\boldsymbol{\omega_{\mathrm{s}}}$, aims to distill knowledge from the main network through the distillation process. Below our definition, the optimization process can be formulated as follows:
$\min_{\omega_{\text{s}}}\mathcal{L}^{\text{d}}_{\text{m}}\big{(}\mathcal{M}(\{\textbf{I}_{\text{vis}},\textbf{I}_{\text{ir}}\};\{\textbf{P}_{\text{vis}},\textbf{P}_{\text{ir}}\};\boldsymbol{\omega_{\mathrm{m}}})\big{)},$ (1)
@@
$s.t.\boldsymbol{\omega_{\mathrm{s}}}^{*}\in\arg\min\limits_{\boldsymbol{\omega_{\mathrm{s}}}}\mathcal{L}^{\mathrm{d}}_{\text{s}}\big{(}\mathcal{S}(\mathbf{I_{\mathrm{vis}}},\mathbf{I_{\mathrm{ir}}};\boldsymbol{\omega_{\mathrm{s}}})\big{)},$ (2)
where, $\mathcal{L}^{\mathrm{d}}_{\text{m}}$ represents the distillation loss for main network, ensuring high-quality fusion while embedding meaningful semantic information, and $\mathcal{L}^{\mathrm{d}}_{\text{s}}$ represents the distillation loss to train the sub-network for effectively mimicking the behavior of the main network. The optimized interaction process is reflected in fused refer image from $\mathcal{M}$, $\mathbf{I_{\mathrm{ref}}}=\mathcal{M}(\{\textbf{I}_{\text{vis}},\textbf{I}_{\text{ir}}\};\{\textbf{P}_{\text{vis}},\textbf{P}_{\text{ir}}\},\omega_{\text{m}})$ and the fused image $\mathbf{I_{\mathrm{f}}}=\mathcal{S}(\mathbf{I_{\mathrm{vis}}},\mathbf{I_{\mathrm{ir}}};\boldsymbol{\omega_{\mathrm{s}}}^{*})$. The goal is to jointly optimize both networks, where the main network learns the rich semantic priors, and the sub-network captures those features under reduced computational requirements.
In the main network, source images are processed by SAM to generate image patches with semantic priors, which, along with features extracted from the source images at multiple scales, are passed through the SPA module to produce fused images. The sub-network serves as a distilled version of the main network. The bi-level optimization distillation scheme, involving triplet losses, facilitates the structured transfer of capabilities from the main network to the sub-network, ensuring that even with reduced complexity. The following sections will detail the design of the SPA module (Section 3.2) and the bi-level optimization distillation scheme involving triplet losses (Section 3.3).
@@ 3.2 Semantic Persistent Attention Module
To fully exploit the semantic information provided by SAM, we propose the Semantic Persistent Attention (SPA) Module, as illustrated in the flowchart in Figure. 3 (b).
@@
The visible and infrared semantic patches ($\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$) extracted from the Semantic Attention Module (SAM) are encoded into features ($\textbf{F}_{\text{pvis}}$ and $\textbf{F}_{\text{pir}}$). These encoded semantic features represent limited portions of the scene and are processed through the cross-attention mechanism. By utilizing the key-value pairs ($\textbf{K}_{\text{src}},\textbf{V}_{\text{src}}$) stored in PR, the cross-attention mechanism enriches the semantic queries ($\textbf{Q}_{\text{pvis}}$ and $\textbf{Q}_{\text{pir}}$) with the full scene context from the source image, thereby addressing the inherent incomplete scene coverage in the patches. PR ensures that the fusion process remains grounded in a stable and consistent context, allowing the patches to be effectively supplemented with modality-specific information from the source. This stable context prevents the fusion process from drifting or losing semantic coherence, while still enabling flexibility in refining the features through attention mechanisms.
@@
Figure 5: Quantitative comparison of fusion performance with other state-of-the-art methods on M3FD and FMB benchmarks. The Violin plots illustrating the distribution of metrics, in which the the black lines and white triangles indicate the medium and mean values.
@@ 3.3 Triplet Loss in Distillation Scheme
In this section, we introduce a triplet loss-driven distillation scheme that facilitates semantic knowledge transfer from the main network to the sub-network, where both networks are optimized in a bidirectional manner. This allows gradients to flow both from the main to the sub-network and vice versa, ensuring a cohesive transfer with mutual compromise. The distillation is driven by three distinct loss components, each targeting different aspects.
The first loss operates at the feature level, aligning the reference image ($\mathbf{I}_{\text{ref}}$) and the fused image ($\mathbf{I}_{\text{f}}$) across multiple scales, which is defined as:
@@
where $\textbf{F}_{\text{Den}}^{m}$ and $\textbf{F}_{\text{SPA}}^{m}$ are the feature maps of the Dense and SPA blocks at the same scale $m$, and $\|\cdot\|_{2}$ is the l2-norm.
@@
where $\nabla$ denotes the Sobel operator and $\|\cdot\|_{1}$ is the l1-norm. So the cont loss can be calculated as $\mathcal{L}_{\text{context}}=\mathcal{L}_{\text{grad}}+\mathcal{L}_{\text{MSE}}$.
@@
$\mathcal{L}_{\text{cs}}^{\text{ir}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{ir}})\right\|_{1}^{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ir}}\odot\mathbf{M}_{\text{ir}})\right\|_{1}^{2}},$ (5)
$\mathcal{L}_{\text{cs}}^{\text{vis}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{vis}})\right\|_{1}^{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{vis}}\odot\mathbf{M}_{\text{vis}})\right\|_{1}^{2}},$ (6)
where $l$ means $l$th layer of $\mathcal{S}_{\text{E}}$, and x $\in\{\text{ref},\text{fus}\}$. So the total $\mathcal{L}_{\text{cs}}=\mathcal{L}_{\text{cs}}^{\text{ir}}+\mathcal{L}_{\text{cs}}^{\text{vis}}$ can also be computed bidirectionally in the bi-level optimization scheme between two networks, ensuring that the feature spaces of both the $\mathbf{I}_{\text{ref}}$ and the $\mathbf{I}_{\text{f}}$ are easily aligned.
The total distillation loss for the sub-network, denoted as $\mathcal{L}^{\text{d}}_{\text{s}}$ in Eq. 2, is the sum of the feature loss, context loss, and contrastive semantic loss: $\mathcal{L}^{\text{d}}_{\text{s}}=\mathcal{L}_{\text{fea}}+\mathcal{L}_{\text{cont}}+\mathcal{L}_{\text{cs}}$.
For the main network, an additional segmentation cross-entropy loss $\mathcal{L}^{\text{seg}}$ is incorporated. This loss is computed between the segmentation predictions generated by an open-vocabulary SAM model $\mathcal{S}^{\text{O}}$ and the ground truth segmentation labels. By including this loss, we aim to leverage the capabilities of SAM to overcome the limitations of traditional segmentation networks, which typically rely on a fixed vocabulary. Instead of affecting the sub-network, this loss aim to prevent potential optimization conflicts. The segmentation loss $\mathcal{L}^{\text{seg}}$ is defined as:
$\mathcal{L}^{\text{seg}}=-\sum_{c}\left[\mathbf{Y}_{\text{label}}^{c}\log(\mathbf{Y}_{\text{pred}}^{c})\right],$ (7)
where $\mathbf{Y}_{\text{label}}$ represents the label segmentation map, $\mathbf{Y}_{\text{pred}}$ is the predicted segmentation map, and $c$ is the class index. For now, the total distillation loss for the main network can be calculated as: $\mathcal{L}^{\text{d}}_{\text{m}}=\mathcal{L}^{\text{d}}_{\text{s}}+\mathcal{L}^{\text{seg}}$.

### v2

@@ Abstract
Multi-modality image fusion, particularly infrared and visible, plays a crucial role in integrating diverse modalities to enhance scene understanding. Although early research prioritized visual quality, preserving fine details and adapting to downstream tasks remains challenging. Recent approaches attempt task-specific design but rarely achieve “The Best of Both Worlds” due to inconsistent optimization goals. To address these issues, we propose a novel method that leverages the semantic knowledge from the Segment Anything Model (SAM) to Grow the quality of fusion results and Enable downstream task adaptability, namely SAGE. Specifically, we design a Semantic Persistent Attention (SPA) Module that efficiently maintains source information via the persistent repository while extracting high-level semantic priors from SAM. More importantly, to eliminate the impractical dependence on SAM during inference, we introduce a bi-level optimization-driven distillation mechanism with triplet losses, which allow the student network to effectively extract knowledge. Extensive experiments show that our method achieves a balance between high-quality visual results and downstream task adaptability while maintaining practical deployment efficiency. The code is available at https://github.com/RollingPlain/SAGE_IVIF.
@@ 1 Introduction
Figure 1: Differences between the proposed method and existing mainstream comparative approaches: (a) Traditional and early DL-based methods focus on the fusion visual effect. (b) Task-specific methods (e.g., TarDAL [11] & SegMiF [13]) introduce task loss and features that lead to inconsistent optimization goals, causing a conflict between visual and task accuracy. (c) Our pipeline first leverages semantic priors from SAM within a large network and then distills the knowledge into a smaller sub-network achieving practical inference feasibility while ensuring “the best of both worlds” through SAM’s inherent adaptability to these tasks.
@@
In light of recent advancements in sensing technologies, multi-modality imaging [15] has gained significant attention, with applications in robotics [35], remote sensing [33], and autonomous driving [1]. Integrating infrared and visible sensors provides substantial advantages for intelligent processing [5], although both exhibit inherent limitations [21, 16]. Infrared images are robust against smoke, obstruction, and low light, while visible images excel in resolution, contrast, and texture detail. These characteristics are highly complementary, motivating their fusion to generate a comprehensive fused image that balances visual quality and downstream task requirements. Consequently, achieving this fusion efficiently remains a pressing challenge [15].
@@
Traditional Infrared and Visible Image Fusion (IVIF) methods, based on information theory [37, 17], aim to retain as much source image information as possible but often struggle with optimizing fused image quality, especially in managing redundancy and specific scenarios. Early deep learning-based methods [32, 8, 12, 40, 39, 22, 18] focused on visualization but faced issues like edge blurring and artifacts, failing to meet downstream perception needs. Current methods [11, 13, 26, 20] focus on specific tasks, coupling fusion with downstream processing, resulting in conflicting optimization objectives and challenges in balancing visual quality with task adaptability.
@@
Recent advancements in large-scale visual models have significantly improved various visual analysis tasks [30]. Among these, the Segment Anything Model [6] stands out due to its exceptional and robust capability to provide rich semantic information, making it well-suited for IVIF, as illustrated in Figure 2. Also, its segmentation capabilities align naturally with the requirements of downstream tasks in IVIF field. However, current methods that integrate SAM for low-level tasks typically require full SAM during inference, resulting in excessive practical infeasibility.
@@
Figure 2: Demonstration of SAM’s robustness in MFNet normal scenes (top) and under challenging FMB conditions (bottom).
@@
To address these issues, we propose the fusion method SAGE that fully integrates and distills semantic priors derived from SAM. As illustrated in Figure 1 III, one of our aim is to harness the advantages of SAM’s semantic priors while reducing computational burden. The proposed method consists of two key components: Semantic Persistent Attention module and a bi-level distillation scheme. The former focuses on using a persistent repository throughout the process to retain source image information, while heuristically enabling the network to integrate semantic patches predicted by SAM. Furthermore, the distillation scheme is designed to efficiently exclude SAM from the inference process, thereby reducing computational complexity. By enhancing distilled information across different aspects, the triplet losses in the scheme significantly boost the performance of the fusion model. Consequently, efficient fusion is achieved using only the distilled model, delivering high-quality visual results and precise task performance without the direct involvement of SAM. In summary, our contributions can be summarized as follows:
  • A novel fusion framework is proposed to leverage semantic priors from the SAM, effectively balancing visual quality and downstream task adaptability..
  • The designed Semantic Persistent Attention (SPA) module utilizes a persistent repository to efficiently retain source information and simultaneously extract high-level semantic representation.
@@
  • We develop a bi-level optimization distillation scheme that transfers the information processed by SPA into the sub-network, effectively decoupling the fusion process from SAM during inference.
@@ 2 Related Works
Learning based IVIF Methods. In recent years, deep learning based multi-modality image fusion approaches [32, 8, 12, 40, 39, 44] achieved significant progress, primarily focusing on improving fusion quality, such as enhancing image details, maintaining structural consistency, and reducing noise. These approaches, including innovations in adversarial learning, Transformer models and diffusion network, have provided new insights. However, most of these methods have not addressed how the fused images can directly contribute to downstream tasks such as object detection, semantic segmentation, and other perception-related applications. With the growing demand for more integrated solutions, this gap has sparked increasing research interest in combining fusion with various downstream tasks.
Early works began to bridge this gap, with TarDAL [11] pioneering the use of a dual-discriminator network for object detection optimization, which simultaneously considers modality differences and detection tasks during fusion. As the field has progressed, several methods [24, 38] have also focused on optimizing fusion for object detection. For semantic segmentation [36], SeAFusion [26] and SegMiF [13] have made early significant improvements, with the former enhancing fusion via task loss and the latter leveraging task-specific features. More recently, some approaches [14, 20, 25, 19, 3, 29] have integrated multi-task learning to jointly optimize detection and segmentation, improving performance on both tasks.
@@
SAM & Its Application.  The Segment Anything Model (SAM) [6] is a large-scale pre-trained foundation model that excels in zero-shot generalization, enabling it to handle a wide range of segmentation and detection tasks without the need for task-specific training. SAM demonstrates impressive performance in low-level vision applications, such as image deblurring [9], dehazing [4], super-resolution [28], and low-light image enhancement [7]. Moreover, SAM’s capabilities extend to high-level tasks like object detection [23], which further strengthens its versatility. This ability to handle both low-level and high-level tasks makes SAM inherently well-suited for the needs of the IVIF field, where fusion based segmentation and detection tasks are critical. However, its impressive performance is hindered by high computational complexity, posing challenges for practical deployment. Addressing these computational burdens, while fully leveraging the semantic priors of SAM, is a primary motivation for our work.
@@ 3.1 Problem Formulation
Figure 4: Qualitative demonstrations of SOTA approaches across commonly used datasets, including TNO, RoadScene, M3FD and FMB.
The overall workflow of our method is illustrated in Figure 3. Our aim is to fully exploit the semantic priors of SAM during the inference stage to enhance cross-modal fusion quality. However, directly deploying a large-scale SAM model often leads to prohibitive computational overhead. To mitigate this, we adopt a knowledge distillation strategy, transferring the information encoded by the SAM-driven main network to a lightweight sub-network, thereby significantly reducing inference costs while maintaining high-quality fusion. Nevertheless, the substantial capacity gap between the SAM-enhanced main network and the compact sub-network frequently results in incomplete semantic transfer or structural inconsistencies, which hinder ideal cross-modal fusion performance. To address this issue, we propose a bi-level optimization framework that jointly optimizes both networks as a unified system, aiming to bridge the distillation gap and maintain consistent fusion guided by SAM’s semantic priors.
@@
Let $\mathbf{I}_{\mathrm{s}}$ denote the source visible and infrared input images, i.e., $\mathbf{I}_{\mathrm{s}}=\{\mathbf{I}_{\mathrm{vis}},\mathbf{I}_{\mathrm{ir}}\}$. Their corresponding semantic-prior patches generated under the guidance of SAM are denoted as $\mathbf{P}_{\mathrm{s}}=\{\mathbf{P}_{\mathrm{vis}},\mathbf{P}_{\mathrm{ir}}\}$. Denote the main network by $\mathcal{M}$ with parameters $\boldsymbol{\omega_{\mathrm{m}}}$, and the sub-network by $\mathcal{S}$ with parameters $\boldsymbol{\omega_{\mathrm{s}}}$. We define the optimization process as:
@@
$\begin{gathered}\min_{\boldsymbol{\omega_{s}}}\mathcal{L}^{\mathrm{d}}_{\mathrm{s}}\big{(}\mathcal{S}(\mathbf{I}_{s},\,\boldsymbol{\omega_{s}});\boldsymbol{\omega_{m}^{*}}\big{)},\\ \text{s.t.}\quad\boldsymbol{\omega_{m}^{*}}\in\arg\min_{\boldsymbol{\omega_{m}}}\mathcal{L}^{\mathrm{d}}_{\mathrm{m}}\big{(}\mathcal{M}(\mathbf{I}_{s},\,\mathbf{P}_{s},\,\boldsymbol{\omega_{m}});\boldsymbol{\omega_{s}}\big{)},\end{gathered}$ (1)
where $\mathcal{L}^{\mathrm{d}}_{\mathrm{m}}$ is the distillation loss that ensures high-quality fusion with meaningful semantic cues in the main network, and $\mathcal{L}^{\mathrm{d}}_{\mathrm{s}}$ is the distillation loss guiding the sub-network to effectively mimic the main network’s behavior. The fused reference image is given by: $\mathbf{I}_{\mathrm{ref}}=\mathcal{M}\big{(}\{\mathbf{I}_{\mathrm{vis}},\mathbf{I}_{\mathrm{ir}}\},\{\mathbf{P}_{\mathrm{vis}},\mathbf{P}_{\mathrm{ir}}\},\boldsymbol{\omega_{m}}\big{)}$, while the sub-network produces a fused image $\mathbf{I}_{\mathrm{f}}=\mathcal{S}\big{(}\mathbf{I}_{\mathrm{vis}},\mathbf{I}_{\mathrm{ir}},\boldsymbol{\omega_{s}}\big{)}$. Notably, both distillation objectives rely on outputs from the counterpart network, establishing a bidirectional dependency, which is a key feature of our distillation formulation.
Within the overall framework, the main network integrates SAM-derived semantic priors into the fusion process, while the sub-network is optimized to align with its outputs under the bi-level formulation. The following sections elaborate on two key components that support this process: the Semantic Persistent Attention (SPA) module 3.2, which integrates SAM-derived priors with source features, and the triplet-based loss scheme 3.3, which facilitates structured knowledge transfer across networks.
@@ 3.2 Semantic Persistent Attention Module
To fully exploit the semantic information provided by SAM, we propose the Semantic Persistent Attention (SPA) module, as illustrated in the flowchart in Figure 3 (b).
@@
The visible and infrared semantic patches ($\textbf{P}_{\text{vis}}$ and $\textbf{P}_{\text{ir}}$) extracted from the SAM are encoded into features ($\textbf{F}_{\text{pvis}}$ and $\textbf{F}_{\text{pir}}$). These encoded semantic features represent limited portions of the scene and are processed through the cross-attention mechanism. By utilizing the key-value pairs ($\textbf{K}_{\text{src}},\textbf{V}_{\text{src}}$) stored in PR, the cross-attention mechanism enriches the semantic queries ($\textbf{Q}_{\text{pvis}}$ and $\textbf{Q}_{\text{pir}}$) with the full scene context from the source image, thereby addressing the inherent incomplete scene coverage in the patches. PR ensures the fusion process stays grounded in a stable, consistent context, enabling patches to be effectively enriched with modality-specific information. This stability maintains semantic coherence while allowing flexible feature refinement via attention mechanisms.
@@
Figure 5: Quantitative comparison of fusion performance with other SOTA methods on M3FD and FMB benchmarks. The violin plots illustrate the distribution of metrics, in which the the black lines and white triangles indicate the medium and mean values.
@@ 3.3 Triplet Loss in Distillation Scheme
In this section, we introduce a triplet loss-driven distillation scheme that facilitates semantic knowledge transfer from the main network ($\mathcal{M}$) to the sub-network ($\mathcal{S}$) under a bi-level optimization framework. Specifically, we adopt a DARTS-style [10] training protocol, where the main network (teacher) and the sub-network (student) are updated in small alternating steps, approximating the inner-outer structure of bilevel optimization. This allows gradients to flow in a bidirectional manner: the sub-network learns from the distillation signals provided by the main network, while the main network also adapts its parameters based on the sub-network’s performance and a segmentation objective. Consequently, the two networks reach a cohesive transfer with mutual compromise.
Our distillation scheme is driven by three distinct losses, each focusing on a different aspect of the fused images. Let $\mathbf{I}_{\text{ref}}$ be the fused reference image produced by the main network, and $\mathbf{I}_{\text{f}}$ be the fused output from the sub-network. We first introduce a feature alignment term:
@@
where $\textbf{F}_{\text{Den}}^{m}$ and $\textbf{F}_{\text{SPA}}^{m}$ are the feature maps of the dense and SPA blocks at the same scale $m$, and $\|\cdot\|_{2}$ is the $\ell_{2}$-norm.
@@
where $\nabla$ denotes the Sobel operator and $\|\cdot\|_{1}$ is the $\ell_{1}$norm. So the cont loss can be calculated as $\mathcal{L}_{\text{context}}=\mathcal{L}_{\text{grad}}+\mathcal{L}_{\text{MSE}}$. In practice, we apply these terms not only between $\mathbf{I}_{\text{ref}}$ and $\mathbf{I}_{\text{fus}}$, but also between each fused output and the original source images to ensure the reconstruction fidelity. This prevents both networks from drifting away from the source image during the distillation process.
@@
$\mathcal{L}_{\text{cs}}^{\text{ir}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{ir}})\right\|_{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{ir}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ir}}\odot\mathbf{M}_{\text{ir}})\right\|_{2}},$ (4)
$\mathcal{L}_{\text{cs}}^{\text{vis}}=\sum_{l=1}^{L}\frac{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{fus}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{ref}}\odot\mathbf{M}_{\text{vis}})\right\|_{2}}{\left\|\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{x}}\odot\mathbf{M}_{\text{vis}})-\mathcal{S}_{\text{E}}(\mathbf{I}_{\text{vis}}\odot\mathbf{M}_{\text{vis}})\right\|_{2}},$ (5)
where $l$ indexes the layers of $\mathcal{S}_{\text{E}}$, and x $\in\{\text{ref},\text{fus}\}$. So the total $\mathcal{L}_{\text{cs}}=\mathcal{L}_{\text{cs}}^{\text{ir}}+\mathcal{L}_{\text{cs}}^{\text{vis}}$ can also be computed bidirectionally in the bi-level optimization scheme between two networks, ensuring that the feature spaces of both the $\mathbf{I}_{\text{ref}}$ and the $\mathbf{I}_{\text{f}}$ are easily aligned.
The total distillation loss for the sub-network, denoted as $\mathcal{L}^{\text{d}}_{\text{s}}$ in Eq. (1), is the sum of the feature loss, context loss, and contrastive semantic loss: $\mathcal{L}^{\text{d}}_{\text{s}}=\mathcal{L}_{\text{fea}}+\mathcal{L}_{\text{cont}}+\mathcal{L}_{\text{cs}}$.
For the main network, an additional segmentation cross-entropy loss $\mathcal{L}_{\text{seg}}$ is incorporated. This loss is computed between the segmentation predictions generated by an open-vocabulary model $\mathcal{S}_{\text{O}}$ [45, 46] and the ground truth segmentation labels. Instead of affecting the sub-network, this loss aim to prevent potential optimization conflicts. The segmentation loss $\mathcal{L}_{\text{seg}}$ is defined as:
$\mathcal{L}_{\text{seg}}=-\sum_{c}\left[\mathbf{Y}_{\text{label}}^{c}\log(\mathbf{Y}_{\text{pred}}^{c})\right],$ (6)
where $\mathbf{Y}_{\text{label}}$ represents the label segmentation map, $\mathbf{Y}_{\text{pred}}$ is the predicted segmentation map, and $c$ is the class index. For now, the total distillation loss for the main network can be calculated as: $\mathcal{L}^{\text{d}}_{\text{m}}=\mathcal{L}^{\text{d}}_{\text{s}}+\mathcal{L}_{\text{seg}}$.